OpenStudy (studybuddy26):

Simplify (2p-py)/-py I solved it myself first and got two possible answers: 3 or 2p+1. Help?!

OpenStudy (steve816):

Welcome to OpenStudy!

OpenStudy (steve816):

\[\Large \frac{ 2p-py }{ -py }\]Is this what you are trying to simplify?

OpenStudy (studybuddy26):

yes

satellite73 (satellite73):

no

satellite73 (satellite73):

\[\Large\frac{ 2-y }{- y }=\frac{ y-2 }{ y }\]

OpenStudy (steve816):

AHhh, oops, simple mistake

OpenStudy (studybuddy26):

Hey found the answer already its 3, and my teach gave me full points for that homework question so don't need help anyone .